# The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages

### v1

@@ Abstract
Developing AI systems that operate effectively across languages while remaining culturally grounded is a long-standing challenge, particularly in low-resource settings. Synthetic data provides a promising avenue, yet its effectiveness in multilingual and multicultural contexts remains underexplored. We investigate the creation and impact of synthetic, culturally contextualized datasets for Indian languages through a bottom-up generation strategy that prompts large open-source LLMs ($\geq 235$B parameters) to ground data generation in language-specific Wikipedia content. This approach complements the dominant top-down paradigm of translating synthetic datasets from high-resource languages such as English. We introduce Updesh111[Uncaptioned image]https://hf.co/datasets/microsoft/Updesh_beta, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages, encompassing diverse reasoning and generative tasks with an emphasis on enhancing long-context and multi-turn capabilities, in addition to improving alignment with Indian cultural contexts. A comprehensive evaluation incorporating both automated metrics and human annotation across 10k assessments indicates that the generated data is of high quality; however, human evaluation highlights specific areas for further improvement. Additionally, we perform downstream evaluations by fine-tuning models on our dataset and assessing their performance across 15 diverse multilingual datasets to evaluate the generalizability of our dataset. Our experiments demonstrate that models trained on Updesh consistently obtain significant performance improvements on generative tasks and remain competitive on multiple-choice evaluations in NLU tasks. Further, the relative improvements of models fine-tuned with Updesh are most pronounced in low and medium-resource languages, effectively narrowing the gap between these languages and high-resource languages. These findings provide empirical evidence that effective multilingual AI development requires multi-faceted data curation and generation strategies that include context-aware, culturally grounded methodologies.
@@ 1 Introduction
Building multilingual and multicultural AI is essential to ensure equitable access to technology in diverse linguistic and cultural communities. Without this inclusivity, AI systems risk reinforcing global power imbalances by privileging dominant languages and Western-centric worldviews. However, even frontier language models often underperform in languages other than English and in non-Western contexts. This is not only due to limited linguistic and cultural diversity in pre-training data, but also to English-centric practices in fine-tuning, evaluation, and other parts of the development pipeline. While pre-training data is often treated as abundant and readily available due to large-scale web crawling, fine-tuning and evaluation data must be deliberately constructed and curated to ensure quality and relevance. This often leads to skipping the creation of truly multilingual and culturally grounded data altogether, or relying on simplified approaches like translating existing English datasets, often overlooking linguistic nuance and cultural context.
As highlighted by Joshi et al. (2020), there is a significant data gap between high-resource and low-resource languages in the internet data, the major source of pre-training data. Their taxonomy classifies world languages into six resource classes, with Classes 5-6 encompassing over $2400$ languages (93.8793.87% of all world languages) serving $1.2$ billion speakers, yet having minimal to no digital resources, creating a massive representation gap in current AI systems. This gap is even more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025). Synthetic data generation has become popular to augment training sets and create benchmarks, by generating new data or by transforming existing data into desirable formats, such as converting raw text into structured instruction-style prompts. Synthetic data, created mainly in English, has shown promise in improving downstream performance in tasks such as reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025) and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025). In this work, we study whether synthetic data has the potential to supplement scarce multilingual, multicultural data when carefully designed.
Techniques for generating synthetic data are typically grounded in English-centric practices, often relying on assumptions about model quality and reliability that do not hold for many other languages. Assessing the quality and utility of synthetic data is an ongoing area of research. Current methods for evaluating quality include human and automated evaluation to assess fluency, correctness, and diversity, which may not suffice for assessing the quality of multilingual, multicultural synthetic data. Downstream evaluation involves fine-tuning models using synthetic data and evaluating models on standardized benchmarks. In this paper, we introduce the first framework for multilingual and multicultural synthetic data generation, focusing on key dimensions such as choice of generation strategies, grounding in language-specific data, quality assessment, and evaluation. We apply this framework to a case study that involves generating synthetic data across 13 Indian languages and conduct comprehensive evaluations to check for quality and utility. Our detailed analysis across multiple multilingual synthetic datasets offers practical insights for developing more inclusive, linguistically diverse, and culturally attuned language technologies.
2 Framework
Synthetic data is utilized throughout the AI lifecycle, including pre-training, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF), and evaluation (Viswanathan et al., 2025). Synthetic data generation starts with a small set of human-curated seeds, which is then provided to an LLM to generate instructions, inputs and outputs, followed by filtering for correctness and quality. An alternate approach is to retrieve the relevant datasets and use an LLM to rewrite them into the desired format. The dominant method for producing multilingual synthetic data involves translating English data with Neural Machine Translation systems or Large Language Models (LLMs) for pre-training (Doshi et al., 2024; Khan et al., 2024), Supervised Fine Tuning (SFT) (Shaham et al., 2024), and evaluation. However, translation may introduce translation artifacts (referred to as translationese) (Zhang & Toral, 2019; Vanmassenhove et al., 2021) and yield weaker correlations with human evaluations compared to benchmarks tailored for particular languages and contexts (Kreutzer et al., 2025; Wu et al., 2025). Recent work has explored bottom-up approaches that rely on grounding and retrieval augmentation, including instruction backtranslation (Li et al., 2024) and instruction back-and-forth translation (Nguyen et al., 2024), which generate high-quality synthetic data grounded in web corpora. These methods can potentially demonstrate superior performance compared to traditional translation-based approaches. However, there remains limited exploration of such approaches in multilingual contexts. In this work, our focus is on curating multilingual, multicultural synthetic data for SFT, although our framework can apply to other kinds of multilingual, multicultural synthetic data, including synthetic pre-training data and preference data as well. Below, we describe the key factors to consider when generating this type of data.
Base model capability: Synthetic data is usually generated by prompting large foundation models due to their fluency, ability to follow instructions, and safety guardrails. For generating multilingual, multicultural synthetic data, candidate base models should be evaluated on established multilingual and multicultural benchmarks to determine performance in target languages. If benchmarks for a specific target language are unavailable, evaluation results from related languages or overall multilingual performance can be used as indicators. Some models list supported languages and may include multilingual benchmark results in their release reports, while others do not provide these details. Additional factors to consider when selecting a base model include licensing, cost, and whether models are open-weights or restricted.
Seed data selection: Seed data is typically selected on the basis of factors such as diversity and target task coverage. Additional considerations for multilingual, multicultural seed data collection would include prioritizing tasks with coverage of cultural knowledge, norms, and values, and tasks with more relevance in specific regional and cultural contexts (for example, agriculture).
Data generation strategy: Several data generation strategies can be considered for synthetic data generation, including translation, back-translation, and retrieval-augmented synthetic data generation. Translating high-quality, curated English SFT datasets allows us to carry over critical skills the LLM needs to acquire into the multilingual training data. Back-translation using existing multilingual datasets can enhance diversity. Retrieval-augmented generation from diverse sources can potentially leverage native speaker-authored content to produce synthetic data that includes cultural knowledge and linguistic nuances.
Quality metrics: Synthetic data should be evaluated for correctness and quality before being used in downstream tasks. This may involve both automated assessment and human evaluation by native speakers. Essential quality metrics for multilingual and multicultural synthetic data include:
  • Language correctness: Is the data in the right language, register, and dialect, or is there confusion with a similar language? (Marchisio et al., 2024)
  • Linguistic acceptability: Does the data sound fluent and natural to a native speaker of the language? (Hada et al., 2024b; a)
  • Cultural appropriateness: Does the data contain appropriate references to cultural artifacts and knowledge and accurately reflect values and norms?
  • Bias and safety: Does the data amplify stereotypes, contain unsafe content, or culturally inappropriate outputs?
Downstream evaluation The primary objective of generating synthetic data is to improve model performance in real-world applications, making the choice of downstream tasks for evaluations critical. This is particularly challenging given the scarcity of multilingual and multicultural benchmarks. To address this, downstream benchmarks should cover all target languages and, wherever possible, avoid relying on translated datasets originally in English. They should also include diverse tasks, especially when synthetic data spans multiple domains. They should incorporate tests of cultural knowledge, norms, and values. In some cases, new benchmarks aligned with the target task, language, and cultural context may need to be developed to accurately assess the utility of synthetic data. Finally, care must be taken to avoid benchmark contamination in the base model, a common issue with multilingual benchmarks (Ahuja et al., 2024).
Native speaker consent, involvement and privacy: Multilingual and multicultural synthetic data creation processes should involve native speakers in seed data selection and evaluation when feasible. Informed consent is required to address cultural considerations and data sovereignty according to local regulations. Synthetic data must not contain personally identifiable information derived from seed data or online sources.
@@ 3 Related Work
English Instruction Fine-tuning Datasets: Instruction Fine-tuning (IFT) (Ouyang et al., 2022) adapts pre-trained language models to follow human instructions through supervised learning on instruction-output pairs. English IFT datasets include the FLAN collection(Wei et al., 2022), which scaled instruction tuning to over 1,800 tasks and chain-of-thought prompting, and Self-Instruct (Wang et al., 2023), which demonstrated effectiveness of LLM-generated synthetic data, leading to Stanford Alpaca (Taori et al., 2023) and the enhanced Alpaca-GPT4 (Peng et al., 2023). Subsequent work expanded through conversational approaches like Vicuna (Chiang et al., 2023) using ShareGPT conversations, WizardLM’s Xu et al. (2024) Evol-Instruct methodology, and LIMA’s (Zhou et al., 2023a) demonstration that 1K curated prompts suffice for effective tuning. The ORCA series (Mukherjee et al., 2023; Mitra et al., 2023) introduced explanation tuning and Prompt Erasure techniques, culminating in OrcaAgent-Instruct (Mitra et al., 2024a) with 25.8M synthetic instruction-response pairs.
@@
Multilingual Instruction Fine-tuning Datasets: Multilingual instruction-following capabilities have been developed through three primary methodologies. Translation-based methods create multilingual datasets by machine translation of existing English instruction data. Bactrian-X (Li et al., 2023) translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) to produce 3.4M pairs across 52 languages, mAlpaca (Chen et al., 2024) demonstrates cost-effective multilingual tuning through Alpaca translation, and LIMA-X achieves 9.9% cross-lingual improvements using $\sim$1K instructions. Template-based approaches like M2Lingual leverage Evol-guided taxonomy (Xu et al., 2024) for synthetic generation across 70 languages. Hybrid strategies combine multiple techniques: Aya-Collection (Singh et al., 2024b) integrates crowdsourcing across 65 languages with templating and translation, repurposing xP3 (Muennighoff et al., 2023), Flan (Longpre et al., 2023), and Dolly datasets while employing NLLB 3.3B (Team et al., 2022) for n-way translation. Similarly, IndicAlign aggregates 74.7M prompt-response pairs across 20 Indian languages through aggregation of existing datasets, translations via the IndicTrans2 (Gala et al., 2023) model, synthetic conversation generation from India-centric Wikipedia, and crowdsourcing.
Data Generation Strategies: Most prior efforts rely on output distillation from advanced LLMs like GPT-4, where models learn from teacher-generated responses. However, recent approaches have explored alternative strategies to address distillation limitations. Instruction Backtranslation (Li et al., 2024) starts with a model fine-tuned on seed data to generate instruction prompts for web documents, followed by self-curation to select high-quality examples. This leverages web content diversity while avoiding over-reliance on teacher models through iterative self-augmentation and self-curation. Instruction Back-and-Forth Translation (Nguyen et al., 2024) extends backtranslation by generating synthetic instructions and then rewriting responses using an LLM to improve quality based on initial documents. This combines web information diversity with high-quality response generation, outperforming distillation and yielding better win rates than ShareGPT and Alpaca-GPT4 (Peng et al., 2023).
Limitations of prior multilingual IFT Datasets: Existing multilingual instruction datasets suffer from critical limitations that hinder effective cross-lingual alignment. Translation-based approaches like Bactrian-X, mAlpaca, and LIMA-X focus primarily on basic instruction-following while excluding advanced reasoning capabilities and lack demographic or cultural representation for local alignment. These datasets suffer from translation inaccuracies due to context-length limitations of traditional sentence-level translation methods (NLLB 3.3B and IndicTrans2) which fail to capture linguistic nuances and introduce errors that can potentially creep in during training. Despite impressive language coverage, Aya-Collection provides limited culturally-specific content, while IndicAlign sources a large proportion of data from WordNet (Miller, 1994) with a focus on question-answering offering limited task diversity. Most existing datasets are short-context and single-turn, failing to address real-world usage patterns including long-context or multi-turn dialogue capabilities. This can significantly constrain model performance in practical applications that demand sustained, contextually-aware interactions.
@@ 4 Data generation
The Updesh dataset comprises two complementary subsets addressing distinct aspects of multilingual instruction following: reasoning data and open-domain generative data. This design recognizes that reasoning capabilities are largely language- and culture-agnostic, making translation-based approaches suitable for tasks like mathematical problem-solving and logical inference (Shaham et al., 2024). Existing high-quality reasoning datasets such as OrcaAgent-Instruct and OrcaMath, thus are valuable resources for multilingual adaptation. However, generative capabilities requiring cultural awareness, linguistic naturalness, and factual grounding in local contexts cannot be adequately addressed through translation due to inherent Western-centric biases and lack of cultural specificity in existing datasets (Yao et al., 2024). Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia-based content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
@@
Reasoning Data: Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages. Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^2] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^3] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table 1). Both datasets have been attributed to induce significant chain-of-thought and reasoning capabilities in models during instruction-tuning without the need for specific preference optimization. We employ Llama-3.1-405B-Instruct for selective translation given its strong coverage in Indian languages and instruction-following capabilities that enable adaptation to various conversational styles (Sankar et al., 2025). Decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and temperature set to $1.0$. After generation, all outputs undergo a series of quality control filters, described in detail below.
Table 1: Reasoning task categories and their descriptions
Task Type Description
Brain Teasers Puzzles stimulating lateral thinking and creativity
Text Classification Categorization tasks for predefined labels
Math Grade-school arithmetic, algebra, and geometry word problems
Open-Domain Generative Data: Synthesizing generative data presents greater challenges than translating existing content, as it introduces increased risks of hallucinations, factual inaccuracies, and demographic misalignment. We conducted a comparative analysis between Llama-3.1-405B-Instruct and Qwen3-235B-A22B models across both reasoning and non-reasoning paradigms to evaluate qualitative differences in generated outputs. Manual assessment revealed that the Qwen3-235B-A22B model demonstrated superior proficiency in generative tasks and excelled in complex instruction following owing to its strong reasoning traces. Therefore, we used the Qwen3-235B-A22B for our generative data curation process. Drawing inspiration from instruction backtranslation techniques (Li et al., 2024), our methodology begins with unlabelled content to construct questions, followed by LLM-generated answers. To ensure generation diversity and contextual grounding that maintains factual accuracy and demographic relevance, we leverage Wikipedia pages in respective target languages as our knowledge base. Table 2 summarizes the eight generative task categories. As demonstrated, some tasks required two phases, which indicates two LLM inferences in order to generate the synthetic data instance.
Task Type Synthesis Method Phases Qwen3-Mode
Further, to ensure cultural representation in synthetic data generation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API. Following (Yao et al., 2024)’s cultural taxonomy framework, we traversed Wikipedia category structures starting from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep. This yielded two complementary datasets: a general collection (54 categories, 19,143 artifacts) covering pan-Indian elements, and a state-wise collection (28 states + 8 union territories, 37,227 artifacts) with region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices. We sampled 26.8K cultural artifacts from this collection to create multi-hop question-answer pairs for synthetic data generation.
Data Filtering After generating approximately 7.5 million data points across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al. (2025), we employed automated quality checks but use the standard threshold-based method instead of their anomaly detection-based method. Specifically, we applied two filtering criteria: (1) Language Identification using IndicLID (Madhani et al., 2023) with a 0.75 confidence threshold, and (2) word repetition ratio capped at 0.75 to flag low-quality generations. Table 3 shows the filtering results, demonstrating high data quality with drop rates below 2% for most subsets. The main exception is the FS-CoT subset for Urdu, where the outputs showed excessive repetition leading to higher filtering rates, but we maintain these thresholds to ensure data integrity. For the English-to-XX translation tasks, Assamese had the highest drop rate as the model frequently generated Bengali text instead, likely due to the shared script and similarity between these languages, and because Assamese is a low-resource language.
Table 3: Document filtering statistics averaged over all Indic Languages
Reasoning Subset Generative Subset
Category Total Drop Rate (%) Category Total Drop Rate (%)
analytical_reasoning 25,000 0.047 logical_reasoning 16,384 1.459
brain_teaser 50,000 0.043 multihop_QA 16,384 1.459
fermi 25,000 0.015 creative_writing 16,384 1.459
fs_cot_flow 25,000 3.769 Multi_turn_Dialog 16,384 1.611
math 200,035 0.035 summarization 16,384 1.526
mcq 99,986 0.135 machine_translation_en 16,384 0.641
rc 50,000 0.379 machine_translation_xx 16,384 17.047
text_classification 50,000 1.878 cultural_multihop_reasoning 26,834 0.347
causal_reasoning 16,384 1.453
5 Data Quality Analysis
5.1 Translation Quality Assessment for the Reasoning Subset
For the reasoning subset, we conducted large-scale selective translation utilizing the Llama-3.1-405B-Instruct. Given that the input often consists of long-context and non-standard text, it is crucial to rigorously evaluate the quality of the generated translations. To this end, following established methodologies the generated output in the target language is translated back into English and compared with the original English source. Due to the substantial volume of data, we randomly sample 4,096 instances from each subset for every language to ensure a representative evaluation. For the backtranslation step, we employed the more efficient Llama-3.3-70B-Instruct, which offers faster inference and serves as a conservative lower bound on translation quality. We assessed translation fidelity using standard metrics such as ChrF (Popović, 2015), implemented using the SACREBLEU library (Post, 2018) for consistency. Table 9 in the Appendix demonstrates consistently high backtranslation scores across all languages and subsets, indicating robust translation quality.
5.2 Data Quality Assessment for Generative Data
Although large language models (LLMs) have emerged as scalable evaluators following the popular LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource language settings still remains limited  (Watts et al., 2024). Our study combined LLM-based assessment with native-speaker human annotation to quantify data quality and establish inter-annotator agreement. We employed a stratified sampling strategy to select 100 data points per task category: Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning tasks (Logical, Causal, and Multi-Hop Reasoning) across five Indian languages (Assamese, Gujarati, Hindi, Malayalam and Punjabi) resulting in a total of 2000 data points. Our sampling methodology ensures representative coverage by partitioning responses into quintiles based on the output length and maintaining proportional sampling across buckets to preserve the original length distribution while capturing variance across responses. Human evaluations were conducted by native speakers through an external agency (Table 7). We employed GPT-4o as an automated evaluator using identical evaluation protocols to allow direct comparison. We devised the following metrics to capture multidimensional quality aspects:
  • Creative Writing: Instruction adherence, linguistic fluency, narrative coherence
  • Reasoning Tasks: Answer adequacy, contextual adherence, instruction compliance, fluency and readability, problematic content detection, cultural relevance assessment (cultural multi-hop only)
  • Multi-Turn Dialogue: Persona consistency, topical coherence, linguistic plausibility, repetition detection, toxicity screening, instruction adherence
  • Summarization: Content coverage, factual accuracy, conciseness, coherence and logical flow, stylistic appropriateness
  • Translation: Semantic correctness, fluency corectness, domain appropriateness, style and tone, completeness
Further details on the metrics and their definitions could be found in the Appendix in Table LABEL:tab:rubrics. All metrics utilize a three-point Likert scale (0-2) with detailed rubrics provided to both human and automated evaluators to ensure assessment consistency. The prompts used for the LLM evaluations could be found in the Appendix A.7. The guidelines given to the human annotators have been added in the supplementary material. Quantitative analysis of human evaluations reveals very high data quality: across 10,000 individual metric assessments, human evaluators assigned only 27 scores of zero (0.27%), indicating that the overwhelming majority of generated content meets acceptable quality thresholds. We plan to publicly release all evaluation data, including both human and GPT assessments, to ensure complete transparency and support future research on calibration of LLM-based evaluators.

### v2

# UPDESH: Synthesizing Grounded Instruction Tuning Data for 13 Indic Languages

@@ Abstract
Developing culturally grounded multilingual AI systems remains challenging, particularly for low-resource languages. While synthetic data offers promise, its effectiveness in multilingual and multicultural contexts is underexplored. We investigate bottom-up synthetic data generation using large open-source LLMs ($\geq 235$B parameters) grounded in language-specific Wikipedia content, complementing dominant top-down translation-based approaches from English. We introduce Updesh, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages and English, encompassing diverse reasoning and generative tasks. Comprehensive evaluation using automated metrics and 10K human assessments confirms high data quality. Downstream evaluations performed by fine-tuning models on various datasets and assessing performance across 13 diverse multilingual datasets and model comparative evaluations, demonstrate that models trained on Updesh consistently obtain significant improvements on NLU, NLG evaluations. Finally, through ablation studies and cultural evaluations, we show that context-aware, culturally grounded data generation is essential for effective multilingual AI development .
UPDESH: Synthesizing Grounded Instruction Tuning Data for 13 Indic Languages
Pranjal A. Chitale   Varun Gummathanks: Work done at Microsoft   Sanchit Ahuja11footnotemark: 1   Prashant Kodali Manan Uppadhyay   Deepthi Sudharsan11footnotemark: 1   Sunayana Sitaramthanks: Corresponding author Microsoft Corporation   Nanyang Technological University   Northeastern University   Independent Researcher pranjalchitale@gmail.com, sunayana.sitaram@microsoft.com
@@ 1 Introduction
Building multilingual, multicultural AI is essential for equitable access across communities. Yet frontier models often underperform in non-English and non-Western settings because diversity is limited in pre-training corpora and English-centric choices pervade the development pipeline (Joshi et al., 2020). While large-scale crawling can expand pre-training data, fine-tuning and evaluation sets require deliberate construction; translation-only approaches often overlook linguistic nuance and cultural context.
Joshi et al. (2020) identify a stark imbalance in web-scale pretraining data: their lowest-resource categories (Classes 5–6) cover over 2.4K languages (93.87% of the world’s languages) spoken by 1.2B people, yet remain severely underrepresented online. This gap is more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025). Synthetic data shows promise in English for reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025), and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025), but pipelines often embed English-centric quality assumptions that may not transfer. Evaluating multilingual, multicultural synthetic data thus remains open: standard human and automatic checks (fluency, correctness, diversity) are often insufficient, and downstream utility must be validated via fine-tuning and benchmarking.
Figure 1: Overview of the data generation pipeline for the Updesh Dataset.
In this work, we introduce Updesh, a culturally grounded multilingual synthetic instruction dataset with 9.5M examples, and provide extensive analyses of its quality and downstream utility. Models fine-tuned on Updesh consistently improve NLG and NLU performance, across languages and shows a considerable uplift in cultural evaluations as well. Alongside the dataset, we distill a set of design considerations for multilingual and multicultural synthetic data generation spanning generation strategies, language-specific grounding, quality assessment, and evaluation. Our results also expose limitations of LLM-as-judge in culturally nuanced settings and yield actionable guidance for dataset design, filtering, and downstream validation, enabling reproducible alternatives to translation-first pipelines.
@@ 2 Related Work
IFT adapts pre-trained LMs to follow instructions using instruction–response pairs (Ouyang et al., 2022). Early English IFT corpora include FLAN (Wei et al., 2022), which scaled to >>1.8k tasks with CoT prompting, and Self-Instruct (Wang et al., 2023), which showed the viability of LLM-generated synthetic data and inspired Stanford Alpaca (Taori et al., 2023) and Alpaca-GPT4 (Peng et al., 2023). Follow-on work explored conversational and curation-heavy routes: Vicuna (Chiang et al., 2023), WizardLM’s Evol-Instruct (Xu et al., 2024), LIMA (Zhou et al., 2023) showing that $\sim$1k high-quality prompts suffice, and the ORCA series (Mukherjee et al., 2023; Mitra et al., 2023), which introduced explanation tuning and prompt erasure, culminating in OrcaAgent-Instruct with 25.8M synthetic pairs (Mitra et al., 2024a).
Multilingual IFT Datasets.
@@
Multilingual instruction-following has been pursued via (i) translation, (ii) template/synthesis, and (iii) hybrids. Translation-focused efforts include Bactrian-X (Li et al., 2023), which translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) into 3.4M pairs over 52 languages, and mAlpaca (Chen et al., 2024), which translates Alpaca. Template-driven generation such as M2Lingual extends Evol-guided taxonomies (Xu et al., 2024) to 70 languages. Hybrid pipelines combine crowdsourcing, templating, and translation: Aya-Collection (Singh et al., 2024b) integrates crowd data across 65 languages with repurposed xP3 (Muennighoff et al., 2023), FLAN (Longpre et al., 2023), and Dolly, using NLLB 3.3B (Team et al., 2022) for N-way translation; IndicAlign aggregates 74.7M prompt–response pairs for 20 Indian languages via dataset aggregation, IndicTrans2-based translation (Gala et al., 2023), synthetic conversations from India-centric Wikipedia, and crowdsourcing.
Data Generation Strategies.
Most prior work distills outputs from stronger teachers (e.g., GPT-4). Recent alternatives mitigate distillation limits by leveraging diverse web content with self-augmentation and self-curation: Instruction Backtranslation (Li et al., 2024) synthesizes instructions from documents, and Back-and-Forth Translation (Nguyen et al., 2024) iteratively rewrites responses with LLMs, often outperforming pure distillation.
Limitations of Prior Multilingual IFT.
Translation-heavy datasets (e.g., Bactrian-X, mAlpaca, and even curation-focused LIMA-X) tend to emphasize basic instruction following, underrepresent advanced reasoning, and provide limited demographic/cultural grounding. Sentence-level MT (e.g., NLLB 3.3B, IndicTrans2) can introduce context-loss and subtle errors that propagate during training. Despite broad coverage, Aya-Collection contains comparatively little culturally specific content, while IndicAlign relies heavily on WordNet (Miller, 1994) and QA-style prompts, limiting task diversity. Finally, most corpora are short-context and single-turn, leaving long-context and multi-turn underexplored.
Reasoning Subset (13 Indian Language & English) Generative Subset (13 Indian Language & English)
Category Total Drop (%) Final Category Total Drop (%) Final
Analytical R 350K 0.047 349.8K Logical R 229.4K 1.459 226.0K
Brain Teaser 700K 0.043 699.7K Multihop QA 229.4K 1.459 226.0K
Fermi 350K 0.015 349.9K Creative Writing 229.4K 1.459 226.0K
Fs-CoT-Flow 350K 3.769 336.8K Multi-turn Dialogue 229.4K 1.611 225.7K
Math 2.80M 0.035 2.80M Summarization 229.4K 1.526 225.9K
MCQ 1.40M 0.135 1.40M Translation (to En) 229.4K 0.641 227.9K
Reading Comp. 700K 0.379 Translation (from En) 229.4K 17.047 190.3K
Text Classification 700K 1.878 686.9K Cultural MHR 375.7K 0.347 374.4K
Causal R 229.4K 1.453 226.0K
Total 7.35M 7.32M Total 2.21M 2.15M
Table 1: Document filtering statistics aggregated over all 13 Indic languages and English. Totals and final counts are reported using K (thousands) and M (millions) notation. Per-language counts are uniform within each category and described in the text. MHR denotes Multi-Hop Reasoning, R denotes Reasoning.
@@ 3 Data Generation
We synthesize Updesh, a dataset covering 13 Indic languages—Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Nepali, Odia, Punjabi, Tamil, Telugu, and Urdu. For each language, Updesh includes two complementary subsets targeting distinct facets of multilingual instruction following: reasoning and open-domain generation. This design recognizes that reasoning capabilities are largely language- and culture-agnostic, making translation-based approaches suitable for tasks like mathematical problem-solving and logical inference (Shaham et al., 2024). We summarize the key design considerations guiding the curation of Updesh in Appendix B.1 along with Figure 6.
@@
Existing high-quality reasoning datasets such as OrcaAgent-Instruct and OrcaMath, thus are valuable resources for multilingual adaptation. However, generative capabilities requiring cultural awareness, linguistic naturalness, and factual grounding in local contexts cannot be adequately addressed through translation due to inherent Western-centric biases and lack of cultural specificity in existing datasets (Yao et al., 2024). Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
Reasoning Data
@@
Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages. Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^1] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^2] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table˜7). Both datasets have been attributed to induce significant chain-of-thought and reasoning capabilities in models during instruction-tuning without the need for specific preference optimization. We employ Llama-3.1-405B-Instruct for selective translation given its strong coverage in Indian languages and instruction-following capabilities that enable adaptation to various conversational styles (Sankar et al., 2025). Post generation, all outputs undergo strict quality checks, as described in detail below.
@@
Open-Domain Generative Data
Synthesizing generative data poses greater challenges than translation due to increased risks of hallucinations, factual inaccuracies, and demographic misalignment. We compared Llama-3.1-405B-Instruct and Qwen3-235B-A22B across reasoning and non-reasoning paradigms, finding Qwen3-235B-A22B superior for generative tasks and complex instruction following due to stronger reasoning traces. (Also supported by Chiang et al. (2024)).
Inspired by instruction backtranslation techniques (Li et al., 2024), we construct questions from unlabelled content followed by LLM-generated answers. To ensure diversity, contextual grounding, factual accuracy, and demographic relevance, we leverage Wikipedia pages in target languages as our knowledge base. Table 8 summarizes eight generative task categories, with some requiring two LLM inference phases. Further, to ensure cultural representation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API.
Following (Yao et al., 2024)’s cultural taxonomy, we traversed Wikipedia categories from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep. This yielded diverse region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices. We sampled 26.8K cultural artifacts to create multi-hop question-answer pairs for synthetic data generation. For English, we reuse the reasoning data from OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) as-is, while generating the English generative subset from scratch using the same pipeline as for the Indian languages.
Data Filtering
After generating data points at scale across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al. (2025), we employed automated quality checks but use the standard threshold-based method instead of their anomaly detection-based method. Specifically, we applied two filtering criteria: (1) Language Identification using IndicLID (Madhani et al., 2023) with a 0.75 confidence threshold, and (2) word repetition ratio capped at 0.75 to flag low-quality generations.
Table 1 shows the filtering results, demonstrating high data quality with drop rates below 2% for most subsets. The main exception is the FS-CoT subset for Urdu, where the outputs showed excessive repetition leading to higher filtering rates, but we maintain these thresholds to ensure data integrity. For the English-to-XX translation tasks, Assamese had the highest drop rate as the model frequently generated Bengali text instead, likely due to the shared script and similarity between these languages, and because Assamese is a low-resource language.
4 Dataset Quality Analysis (Q-A)
4.1 Q-A for Reasoning Data
For the reasoning subset, we performed large-scale selective translation using Llama-3.1-405B-Instruct. Given inputs with long contexts and non-standard text, we rigorously evaluated translation quality through backtranslation. We randomly selected 4,096 samples per subset and language, backtranslated them to English using Llama-3.3-70B-Instruct (chosen for faster inference and conservative quality bounds), and compared with original sources. Translation fidelity was measured using ChrF (Popović, 2015) via SACREBLEU (Post, 2018). Table 10 shows consistently high backtranslation scores across all languages, confirming robust translation quality.
4.2 Q-A for Generative Data
While large language models (LLMs) have become scalable evaluators under the LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource settings remains limited (Watts et al., 2024; Whitehouse et al., 2025). We therefore combined LLM evaluation with native-speaker annotation and measured inter-annotator agreement. Using stratified sampling, we drew 100 instances per category—Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning (Logical, Causal, MultiHop)—across five languages (Assamese, Gujarati, Hindi, Malayalam, Punjabi), yielding 2K items. Sampling preserved response-length distributions via quintile bucketing. Native-speaker evaluations were conducted via an external agency (Table 6), while GPT-4o served as an automated evaluator using identical protocols for comparability. We defined task-specific, multidimensional metrics using a 3-point Likert scale (0–2) with consistent rubrics (Section B.3), ensuring a thorough quality check of the generated data. Prompts and annotation guidelines are provided in Appendix C.1 and supplementary material. Across 10K individual ratings, only 27 received a zero (0.27%), indicating uniformly high data quality. We will publicly release all evaluation data—including both human and GPT assessments—to promote research on calibrating LLM-based evaluators.